# Understanding User Mental Models in AI-Driven Code Completion Tools: Insights from an Elicitation Study

### v1

@@ 4.2 Explanations of the completion
The first aspect we analyzed regarding suggested code explanations concerns where the suggestion should be displayed in the user interface of the CCTs. The first strategy concerned using a separate window or pop-up (7 groups). Several participants proposed that the explanations are better suited to a separate interface element, such as a sidebar or pop-up, allowing them to access the explanations without cluttering up the coding work area (“The explanation and the function must be shown separately, otherwise there’s too much text” – G1). Another strategy, on the other hand, concerns visualization employing inline comments within the code (6 groups), thus providing an immediate context for the suggested code without having to navigate elsewhere (“As an expandible text under the code” – G7); however, this solution was indicated as viable only if the explanation length was limited, in order to avoid confusion in the code (“No comments, as they tend to be too long… but a ‘title’ of the suggestion can be kept as a brief explanation in a comment” – G6). The last solution concerns using chatbots (3 groups). Some participants suggested accessing the explanations via a chatbot, allowing users to query the system in a conversational manner to obtain further clarification of the suggestions (“An alternative could be the chatbot to which I ask why [the code was suggested] after selecting the code” – G8).
@@
Three complementary strategies emerged concerning the timing of explanations, i.e., when explanations should be shown during the interaction with the CCTs. The first solution is to activate the explanations by manually interacting with the suggested code (6 groups). For example, users can select the piece of suggested code they want to explain, right-click, then choose an ad-hoc function from a contextual menu, and the explanation is then displayed in a separate window, in the chatbot, or as a comment (“For example with the right click on the code and then pressing ‘explain’” – G1); hovering over suggested code to show explanations contextually was also proposed by some groups (“… and then with an explicit request … hovering on the code” – G6). An alternative solution concerns activation via shortcuts (3 groups). Indeed, some participants suggested activating the explanations through a dedicated keyboard shortcut, granting users control without interfering with their workflow (“With a key I invoke the explanation” – G8). Finally, an automatic approach along with the suggestion has also been proposed (3 groups), i.e., also accompanying the suggested code with an automatic short explanation, expandable on-demand, which is especially useful when the suggestion is complex, reducing cognitive effort by presenting both simultaneously. A longer suggestion should be provided on-demand, for example, by interacting with the short suggestion and invoking more details (“… generated together with the code, as long as it is brief” – G2; “[explanation] details are provided on-demand” – G3).
The last aspect concerns what to explain. The desired content of the explanations was varied, with participants identifying three key elements. The first aspect concerns purpose (5 groups), i.e., explaining why a certain code suggestion was given. Indeed, many participants emphasized the need for explanations that clarify the purpose of the suggested code, specifying what it achieves in the broader context of the program (“[explain] why did it give me this suggestion?” – G6; “[explain] pros and cons of a suggested technique” – G8). The second main aspect regards functionality (4 groups), i.e., describing how the suggested code fulfills its purpose by reporting technical details such as the data structures, code constructs, data types and syntax involved (“What is the code composed of (in terms of data structures and chosen constructs)?” – G6; “explain the code, data types, … syntax” – G3); some groups also suggested displaying flow diagrams or pseudo-code to help developers understand complex suggestions and visualize the inner workings of the generated code (“With pseudocode, maybe also flowcharts” – G1). Another popular aspect concerns the use of examples and references to sources (5 groups): participants recommended that explanations include examples of the use of similar codes or references to reliable sources that might enable users to assess the credibility of the suggestion (“It may be sufficient to provide simple ‘pointers’ to something that helps me understand … such as links to source codes, or similar code examples as ‘insights’” – G2). Regarding the scope of explanations, it emerged that explanations should be contextual (4 groups) and not regard the entirety of the suggested code, but only the pieces of code that the user indicates (“We’re not interested in being explained everything necessarily, but only of a piece [of code]” – G1). A final aspect concerns the length of explanations: there was a popular agreement that explanations should be minimal and provide more details on request (4 groups). Participants suggested that this can be achieved by, e.g., clicking a button to expand the explanation preview or by interacting with a chatbot to obtain an explanation incrementally (“If something else specific is needed, the user can ask explicitly” – G3).
@@ 5 Discussion
Our findings highlight important points regarding developer preferences and requirements while working with AI-driven CCTs. In particular, this work underlines the importance of designing CCTs for developers’ mental models to promote productivity, efficiency, and an intuitive coding experience. In the discussion reported in the following sub-sections, the interpretation of these results reveals how they challenge our current CCT design paradigms and point toward improvements in CCT design. Figure 3 summarizes users’ preferences and requirements, attempting to elicit developers’ mental model of CCTs.
@@
5.1 Activation and control
@@
Participants expressed different preferences regarding when code suggestions by CCT should be triggered, namely proactive, manual, and hybrid activation. This diversity in the mental model points to a key insight: in different contexts and coding tasks, developers might prefer different degrees of control. Some users mentioned that proactive suggestions, automatically activated during the coding flow breaks, might increase coding speed without explicit request. However, most users expressed the need for activation on-demand to control exactly when suggestions appeared. This is in line with similar studies regarding activation, as frequently CCTs are not adopted as they might disrupt developers’ workflow [2]. Rather, users favored a hybrid approach in which proactive behavior occurs only when a clear contextual need is present (such as when completing a complicated bit of code). However, proactive behavior should only occur once there is enough typing context (e.g., suggest code only if the file contains X tokens or more) to avoid generating unprecise or random suggestions. Therefore, a dynamic and adaptive activation model could serve different user needs by altering its behavior in response to task complexity and user-characteristic usage patterns. Such a model would leave developers in control of their workflow while helping proactively when needed.
5.2 Visualization of suggestions
@@
The visualization mechanism of suggestions is crucial to the usability and acceptance of CCTs. The participants proposed different visualization strategies, each with pros and cons. The proposed strategies are inline suggestions shown with a different font, suggestions displayed in sidebars, or suggestions visualized in pop-ups. Having inline suggestions can be useful to avoid losing the coding focus and allow the user to accept or reject the suggestion immediately. However, the inline solution is well suited for short and concise suggestions that do not interfere too much with the code manually written by the users; on the contrary, in the case of long and complex suggestions, it may be better to have separate windows or pop-ups, to limit visual clutter. A further level of control would be having users manually invoke popups containing the suggestion by clicking an icon shown close to the handwritten code. Moreover, a chatbot in a separate window could be used as an alternative way to suggest code on-demand, with all the attached benefits of using conversational agents for code generation [44].
@@
All these results may reveal an adaptable visualization, where users can choose or alternate between inline, sidebar, and popups, and an adaptive visualization, where the system visualizes the suggestion according to factors such as the suggestion length. Such modalities can improve usability by adapting to different workflows and developer preferences and might reduce the errors due to wrong intention communication as multiple choices may be provided when needed [2]. This flexibility can also lower the cognitive overhead caused, for instance, by long pieces of generated code shown next to user-written code; in this way, CCT allows developers to adjust the display of suggestions according to the complexity of the coding task or their personal coding style.
5.3 Content and granularity
@@
The study findings revealed that CCTs should be capable of generating suggestions at varying levels of granularity and that this granularity must be in the control of the users. Participants did not clearly agree on the minimum or maximum amount of tokens that should be suggested at any given time. Consequently, the CCTs should be designed to produce suggestions at different scopes, from predicting the next token in the code to completing control structures and generating entire functions or classes.
@@
Although longer suggestions are useful, they can overwhelm the developer if displayed without considering the task context. For this reason, participants prefer retaining control over the level of granularity of suggestions. The necessity for varying degrees of granularity is also shared by other participants in different studies [8]. However, in general, users expressed a preference towards minimal, “lower-scope”, suggestions at the first moment, with the possibility to obtain a longer, “higher-scope”, suggestion on-demand; this approach of “letting the user reason” could help to prevent overreliance and deskilling. In practice, these directives might translate concretely into a minimalist suggestions behavior by default, with the possibility of using a slider or a control mechanism to adjust granularity in real time and adapt the suggestions to the contextual needs of the developer.
@@ 5.4 Explanation preferences
Participants generally considered explanations useful, but the answers to the question of what should be contained in an explanation revealed a need for special care in their design. The concept of what defines a good explanation is not trivial and is an aspect that is widely studied in the literature [45, 46]. Moreover, even well-designed explanations should be evaluated with users to measure their quality [46]. This elicitation study can be a precious source of information in understanding the desiderata of explanations in CCTs from the developers’ point of view. A key finding is that explanations should be classified into two categories: why-explanations and what-explanations. Why-explanations address the rationale or purpose behind the generated code, whereas what-explanations concentrate on the functionality and technical intricacies of the code suggestion. The content should, however, be minimal, with the possibility to obtain more details when requested. This might translate into the CCT by initially showing a brief why-explanation, providing a bird-eye view of what the generated code is for. A more thorough, on-demand explanation would instead detail the code nuances and technical aspects to help the developer get a deeper understanding of the suggestion, possibly even with flowcharts. Detailed explanations should also include references to trusted sources, such as documentation snippets, book pieces, etc., and even examples of use to enhance understanding.
@@
Regarding where to show explanations, participants typically preferred separate elements (e.g., popups, sidebars), mainly because they did not want the explanations to be invasive and occupy space within the coding area. However, inline explanations may be accepted as comments close to the suggested code as long as their length is very limited. However, separate windows may allow developers to initiate an explanation process with a conversational agent, similar to what is possible to achieve with tools such as ChatGPT [47].
@@
Timing also plays a crucial role in showing the explanation. Some participants stated that explanations should be triggered manually using shortcuts, mouse clicks, or mouse hovering; notably, selecting code to explain with the mouse was a popular preference, also because it allows limiting the explanation to the context of the selection. Such an approach would have the advantage of reducing the amount of non-relevant explanation text to show to the user; moreover, it would also allow asking for further detailed explanations limited to a specific piece of manually selected code, possibly to a chatbot. Some other participants instead expressed that brief explanations should be shown automatically (especially for novice developers), letting the user choose whether to receive more details on demand. Thus, all these findings indicate that adaptive, context-aware explanations can increase the understanding and trust of a tool dealing with complex or ambiguous code suggestions.
5.5 Personalization needs
@@
An aspect that emerged as essential to fit the CCT’s behavior in different mental models is the explicit personalization of the CCT. There was indeed little consensus among participants towards one universally accepted solution; in fact, many individual factors, such as experience with coding, quality requirements, and personal preferences, play a role in determining the optimal behavior (and, therefore, acceptance) of the CCT. For example, developers want to customize when suggestions should appear, how to visualize them, and which granularity should be used by default. Choosing whether suggestions should also be suggested automatically resulted in being, indeed, crucial to users for the acceptance of the CCT, while they also expressed the preference to edit what manual action (e.g., shortcut, toggle button, etc.) activates the code completion. In addition, the possibility of customizing the suggestions’ font, colors, and size and controlling the suggestions’ position within the IDE also proved useful. Another critical point regarded the need to customize the granularity of suggestions to control the amount of code provided by the tool, both regarding the scope of the suggestion (function-level, control structure-level, etc.) and the number of alternatives provided.
@@
Moreover, users expressed the need to customize aspects regarding generating explanations regarding the suggested code. Therefore, it is important for developers to customize aspects such as the activation timing of explanations (whether to provide them proactively) and their granularity (such as the amount and frequency of comments within the generated code and the explanation abstraction).
@@
6 Prototype Development
The insights that emerged during the elicitation study represent a set of actionable guidelines for CCTs designers. To verify and demonstrate the feasibility of these guidelines, we developed ATHENA (AI Tool for Human-centered ENhanced coding Assistance), a proof of concept of a CCT for the Visual Studio Code IDE. The tool’s name was inspired by Athena, the Greek goddess of arts, crafts, and wisdom, and symbolizes intelligence, skill, and creativity—all attributes that align well with a tool designed to support developers. Developers can view the tool as a source of guidance, precision, and mastery, like Athena represents strategic thinking and expertise. The tool is available as open source software at the following link: https://github.com/IVU-Laboratory/ATHENA.

### v2

@@ 5.2 Explanations of the completion
The results of this subsection answer RQ2, which investigates developers’ expectations on how explanations should be delivered by CCTs. The study findings clarify user preferences about the timing, format, and content of explanations. In addition, the need for a customizable level of detail and contextual delivery answers RQ3, highlighting the importance of tailoring explanations to user expertise and task demands. Figure˜3 represents the identified themes.
@@
Where to show the explanation in the UI
The first aspect we analyzed regarding suggested code explanations concerns where the explanations should be displayed in the user interface of the CCTs. Three suggestions emerged:
  • Separate window or pop-up (7 groups): several participants proposed that the explanations are better suited to a separate interface element, such as a sidebar or pop-up, allowing them to access the explanations without cluttering up the coding work area (“The explanation and the function must be shown separately, otherwise there’s too much text” – G1).
@@
  • Inline comments (6 groups): another strategy concerns visualization employing inline comments within the code, thus providing an immediate context for the suggested code without having to navigate elsewhere (“As an expandible text under the code” – G7); however, this solution was indicated as viable only if the explanation length was limited, in order to avoid confusion in the code (“No comments, as they tend to be too long… but a ‘title’ of the suggestion can be kept as a brief explanation in a comment” – G6).
  • Chatbots (3 groups): some participants suggested accessing the explanations via a chatbot, allowing users to query the system in a conversational manner to obtain further clarification of the suggestions (“An alternative could be the chatbot to which I ask why [the code was suggested] after selecting the code” – G8).
When to show the explanation during the interaction
Three complementary strategies emerged concerning the timing of explanations, i.e., when explanations should be shown during the interaction with the CCTs:
@@
  • Manually interacting with the suggested code (6 groups): the first solution is to activate the explanations by manually interacting with the suggested code. For example, users can select the piece of suggested code they want to explain, right-click, then choose an ad-hoc function from a contextual menu, and the explanation is then displayed in a separate window, in the chatbot, or as a comment (“For example with the right click on the code and then pressing ‘explain’” – G1); hovering over suggested code to show explanations contextually was also proposed by some groups (“… and then with an explicit request … hovering on the code” – G6).
  • Shortcut (3 groups): an alternative solution concerns activation via shortcuts. Indeed, some participants suggested activating the explanations through a dedicated keyboard shortcut, granting users control without interfering with their workflow (“With a key I invoke the explanation” – G8).
  • Accompany the suggested code (3 groups): an automatic approach along with the suggestion has also been proposed, i.e., also accompanying the suggested code with an automatic short explanation, expandable on-demand, which is especially useful when the suggestion is complex, reducing cognitive effort by presenting both simultaneously. A longer suggestion should be provided on-demand, for example, by interacting with the short suggestion and invoking more details (“… generated together with the code, as long as it is brief” – G2; “[explanation] details are provided on-demand” – G3).
What should be explained
The last aspect concerns what to explain. The desired content of the explanations was varied, with participants identifying three key elements:
  • Purpose (5 groups): explain why a certain code suggestion was given. Indeed, many participants emphasized the need for explanations that clarify the purpose of the suggested code, specifying what it achieves in the broader context of the program (“[explain] why did it give me this suggestion?” – G6; “[explain] pros and cons of a suggested technique” – G8).
  • Functionality (4 groups): describe how the suggested code fulfills its purpose by reporting technical details such as the data structures, code constructs, data types and syntax involved (“What is the code composed of (in terms of data structures and chosen constructs)?” – G6; “explain the code, data types, … syntax” – G3); some groups also suggested displaying flow diagrams or pseudo-code to help developers understand complex suggestions and visualize the inner workings of the generated code (“With pseudocode, maybe also flowcharts” – G1).
  • Examples and references (5 groups): another popular aspect concerns the use of examples and references to sources : participants recommended that explanations include examples of the use of similar codes or references to reliable sources that might enable users to assess the credibility of the suggestion (“It may be sufficient to provide simple ‘pointers’ to something that helps me understand … such as links to source codes, or similar code examples as ‘insights’” – G2).
  • Explanations should be contextual (4 groups): regarding the scope of explanations, they should be contextual and not regard the entirety of the suggested code, but only the pieces of code that the user indicates (“We’re not interested in being explained everything necessarily, but only of a piece [of code]” – G1).
  • Explanations should be minimal and provide more details on request (4 groups): regarding the length of explanations, there was a popular agreement that explanations should be minimal and provide more details on request. Participants suggested that this can be achieved by, e.g., clicking a button to expand the explanation preview or by interacting with a chatbot to obtain an explanation incrementally (“If something else specific is needed, the user can ask explicitly” – G3).
@@ 6 Discussion
Our findings highlight important points regarding developer preferences and requirements while working with AI-driven CCTs. In particular, this work underlines the importance of designing CCTs for developers’ mental models to promote productivity, efficiency, and an intuitive coding experience. Figure˜4 summarizes users’ preferences and requirements, attempting to elicit developers’ mental model of CCTs.
This section interprets and presents the findings of our study in light of the research questions. Specifically, Sections˜5.1 and 5.2 addressed RQ1, RQ2, and RQ3 by identifying how developers mentally model their interactions with AI-driven CCTs, what they expect from explanations, and how they express their need for personalization. More broadly, this section addresses RQ4 by distilling the study results into actionable human-centered design suggestions, which are intended to guide the development of CCTs that align with developers’ mental models and support diverse coding workflows.
@@
6.1 Mental model of activation and control (RQ1)
Participants expressed different preferences regarding when code suggestions by CCT should be triggered, namely proactive, manual, and hybrid activation. This aspect is supported by the results detailed in Section˜5.1 (‘When to show suggestions?’) and Table˜3, where proactive activation was favored by 4 groups, manual by 5 groups, and hybrid by 6 groups.
@@ 6.1 Mental model of activation and control (RQ1)
This diversity in the mental model points to a key insight: in different contexts and coding tasks, developers might prefer different degrees of control. Some users mentioned that proactive suggestions, automatically activated during the coding flow breaks, might increase coding speed without explicit request. However, most users expressed the need for activation on-demand to control exactly when suggestions appeared. This is in line with similar studies regarding activation, as frequently CCTs are not adopted as they might disrupt developers’ workflow (Sergeyuk et al., 2025). Rather, users favored a hybrid approach in which proactive behavior occurs only when a clear contextual need is present (such as when completing a complicated bit of code). However, proactive behavior should only occur once there is enough typing context (e.g., suggest code only if the file contains $X$ tokens or more) to avoid generating imprecise or random suggestions. Therefore, a dynamic and adaptive activation model could serve different user needs by altering its behavior in response to task complexity and user-characteristic usage patterns. Such a model would leave developers in control of their workflow while helping proactively when needed.
6.2 Visualization preferences (RQ1)
@@
The visualization mechanism of suggestions impacts the usability and acceptance of CCTs. The participants proposed different visualization strategies, each with pros and cons. As shown in Section˜5.1 (‘How to show suggestions?’) and summarized in Fig.˜2, the majority of groups (5/8) preferred inline suggestions, as well as suggestions displayed in sidebars (7/8). Having inline suggestions can be useful to avoid losing the coding focus and allow the user to accept or reject the suggestion immediately. However, the inline solution is well-suited for short and concise suggestions that do not interfere too much with the code manually written by the users; on the contrary, in the case of long and complex suggestions, it may be better to have separate sidebars to limit visual clutter. A further level of control would be having users manually invoke popups containing the suggestion by clicking an icon shown close to the handwritten code (proposed by 2 groups). Moreover, a chatbot in a separate window could be used as an alternative way to suggest code on demand (proposed by 2 groups), with all the attached benefits of using conversational agents for code generation (Ross et al., 2023).
@@
All these results may reveal an adaptable visualization, where users can choose or alternate between inline, sidebar, and popups, and an adaptive visualization, where the system visualizes the suggestion according to factors such as the suggestion length. Such modalities can improve usability by adapting to different workflows and developer preferences and might reduce the errors due to wrong intention communication as multiple choices may be provided when needed (Sergeyuk et al., 2025). This flexibility can also lower the cognitive overhead caused, for instance, by long pieces of generated code shown next to user-written code; in this way, CCT allows developers to adjust the display of suggestions according to the complexity of the coding task or their personal coding style.
6.3 Granularity and content of suggestions (RQ1, RQ3)
The study findings revealed that CCTs should be capable of generating suggestions at varying levels of granularity and that this granularity must be in the control of the users. This finding traces to Section˜5.1 and Table˜5, where it is reported that 7 out of 8 groups supported user control over suggestion granularity, including requests for minimal suggestions with additional detail on-demand found in 4 groups.
@@ 6.3 Granularity and content of suggestions (RQ1, RQ3)
Participants did not clearly agree on the minimum or maximum number of tokens that should be suggested at any given time. Consequently, the CCTs should be designed to produce suggestions at different scopes, from predicting the next token in the code to completing control structures and generating entire functions or classes.
@@
Although longer suggestions are useful, they can overwhelm the developer if displayed without considering the task context. For this reason, participants prefer retaining control over the level of granularity of suggestions. The necessity for varying degrees of granularity is also shared by other participants in different studies (Wang et al., 2023). However, in general, users expressed a preference towards minimal, “lower-scope”, suggestions at the first moment, with the possibility to obtain a longer, “higher-scope”, suggestion on-demand; this approach of “letting the user reason” could help to prevent overreliance and deskilling. In practice, these directives might translate concretely into a minimalist suggestions behavior by default, with the possibility of using a slider or a control mechanism to adjust granularity in real time and adapt the suggestions to the contextual needs of the developer.
@@ 6.4 Preferences for explanation delivery (RQ2)
Participants generally considered explanations useful, but the answers to the question of what should be contained in an explanation revealed a need for special care in their design. The concept of what defines a good explanation is not trivial and is an aspect that is widely studied in the literature (Miller, 2019; Holzinger et al., 2020). Moreover, even well-designed explanations should be evaluated with users to measure their quality (Holzinger et al., 2020). This elicitation study can be a precious source of information in understanding the desiderata of explanations in CCTs from the developers’ point of view. A key finding is that explanations should be classified into two categories: why-explanations and what-explanations. Why-explanations address the rationale or purpose behind the generated code, whereas what-explanations concentrate on the functionality and technical intricacies of the code suggestion. The content should, however, be minimal, with the possibility to obtain more details when requested. This might translate into the CCT by initially showing a brief why-explanation, providing a bird-eye view of what the generated code is for. A more thorough, on-demand explanation would instead detail the code nuances and technical aspects to help the developer get a deeper understanding of the suggestion, possibly even with flowcharts. Detailed explanations should also include references to trusted sources, such as documentation snippets, book pieces, etc., and even examples of use to enhance understanding.
@@
Regarding where to show explanations, participants typically preferred separate elements (e.g., popups, sidebars), mainly because they did not want the explanations to be invasive and occupy space within the coding area. However, inline explanations may be accepted as comments close to the suggested code as long as their length is very limited. However, separate windows may allow developers to initiate an explanation process with a conversational agent, similar to what is possible to achieve with tools such as ChatGPT (OpenAI, 2024a).
@@
Timing also plays a crucial role in showing the explanation. Some participants stated that explanations should be triggered manually using shortcuts, mouse clicks, or mouse hovering; notably, selecting code to explain with the mouse was a popular preference, also because it allows limiting the explanation to the context of the selection. Such an approach would have the advantage of reducing the amount of non-relevant explanation text to show to the user; moreover, it would also allow asking for further detailed explanations, limited to a specific piece of manually selected code, possibly to a chatbot. Some other participants instead expressed that brief explanations should be shown automatically (especially for novice developers), letting the user choose whether to receive more details on demand. Thus, all these findings indicate that adaptive, context-aware explanations can increase the understanding and trust of a tool dealing with complex or ambiguous code suggestions.
Participant preferences for delivering explanations are evidenced in Section˜5.2, Table˜7 (7/8 groups favoring a separate window or pop-up), Table˜8 (6/8 groups preferring manual interaction), and Table˜9 (4/8 groups requesting minimal explanation with more details on request).
6.5 Personalization and design needs (RQ3, RQ4)
@@
An aspect that emerged as essential to fit the CCT’s behavior in different mental models is the explicit personalization of the CCT. There was, indeed, little consensus among participants towards one universally accepted solution; in fact, many individual factors, such as experience with coding, quality requirements, and personal preferences, play a role in determining the optimal behavior (and, therefore, acceptance) of the CCT. For example, developers want to customize when suggestions should appear, how to visualize them, and which granularity should be used by default. Choosing whether suggestions should also be suggested automatically resulted in being, indeed, crucial to users for the acceptance of the CCT, while they also expressed the preference to edit what manual action (e.g., shortcut, toggle button, etc.) activates the code completion. In addition, customizing the suggestions’ font, colors, and size and controlling the suggestions’ position within the IDE also proved useful. Another critical point regarded the need to customize the granularity of suggestions to control the amount of code provided by the tool, both regarding the scope of the suggestion (function-level, control structure-level, etc.) and the number of provided alternatives.
@@
The users expressed the need to customize aspects of generating explanations regarding the suggested code. Therefore, developers must be able to customize aspects such as the activation timing of explanations (whether to provide them proactively) and their granularity (such as the amount and frequency of comments within the generated code and the explanation abstraction).
@@
7 Reifying the mental model: the ATHENA prototype
In this section, we put the mental model derived from our elicitation study into practice by presenting ATHENA, a prototype implementation that embodies and operationalizes the key design principles and the identified user interaction preferences.
ATHENA (AI Tool for Human-centered ENhanced coding Assistance) is a proof-of-concept CCT for Visual Studio Code IDE, developed to verify and demonstrate the feasibility of the actionable guidelines emerging from our study. The implementation incorporates high configurability, such as suggestion timing, granularity control, adaptive visualization (inline, sidebar, popup), and customizable explanation delivery that allows developers to tailor the extension to their preferences and workflows. As its namesake, Athena, the Greek goddess of skill and strategy, the tool aims to provide developers with precise, adaptable coding support.